# GeoReasoner: Geo-localization with Reasoning in Street Views using a Large Vision-Language Model

### v1

@@ 2.2 Image-based Geo-localization
We approach image-based geo-localization with a novel paradigm. Specifically, we integrate semantic visual concepts that offer locatable features (Theiner et al., 2022), and incorporate human reasoning knowledge learned from geo-localization games using an LVLM.
@@ 3.1 Locatability-Enhanced Data Curation
  • Textual Clues. Textual clues often serve a pivotal role in delineating the geographical locations of street-view images. However, there is currently no readily available textual clue dataset. Two prominent games, namely GeoGuessr [^3] https://www.geoguessr.com and Tuxun [^4] https://tuxun.fun, which are centered around geo-localization based on street views, provide a potential solution to this gap. The communities surrounding these games possess a well-structured and efficacious collection of textual clues, utilized for geographical positioning across various countries and cities. These clues are collaboratively maintained and curated by both administrators and players. The domain knowledge lays the groundwork for identifying decisive factors and assessing the significance of each factor in street views. We acquired image-text data pairs for geo-localization from the two open-source communities. Subsequently, we utilized the BERT-based Named Entity Recognition (NER) (Kenton & Toutanova, 2019) model to clean and filter text that lacked specific geographical location information. In this way, we collected a total of over 3K textual clues that encapsulate rich geo-localization information. For instance, “houses in central Chile are more likely to have terracotta tiled roofs”. Each clue is paired with a corresponding street-view image.

### v2

@@ 2.2 Image-based Geo-localization
We approach image-based geo-localization with a novel paradigm. Specifically, we integrate semantic visual concepts that offer locatable features (Luo et al., 2022; Theiner et al., 2022), and incorporate human reasoning knowledge learned from geo-localization games using an LVLM.
@@ 3.1 Locatability-Enhanced Data Curation
  • Textual Clues. Textual clues often serve a pivotal role in delineating the geographical locations of street-view images. Two prominent games, GeoGuessr [^3] https://www.geoguessr.com and Tuxun [^4] https://tuxun.fun, which focus on geo-localization through street views, offer a potential solution to this gap. Their communities have collaboratively curated a well-organized collection of textual clues, used for pinpointing geographical locations across various countries and cities. These clues, maintained by both players and administrators, provide valuable domain knowledge that aids in identifying and evaluating key geographical features in street views. While such datasets now exist (Luo et al., 2022), there are no readily available image-text data pairs specifically tailored for LVLM training. To bridge this gap, we gathered image-text pairs for geo-localization from these two open-source communities. Subsequently, we utilized the BERT-based Named Entity Recognition (NER) (Kenton & Toutanova, 2019) model to clean and filter text that lacked specific geographical location information. In this way, we collected a total of over 3K textual clues that encapsulate rich geo-localization information. For instance, “houses in central Chile are more likely to have terracotta tiled roofs”. Each clue is paired with a corresponding street-view image.